The following table shows a data set and the number of times each point is misclassified during a run of the perceptron algorithm (with offset θ0). θ and θ0 are initialized to zero.

i x(i) y(i) times misclassified
1 [-4, 2] +1 1
2 [-2, 1] +1 0
3 [-1, -1] -1 2
4 [2, 2] -1 1
5 [1, -2] -1 0

Write down the state of θ and θ0 after this run has completed (note, the algorithm may not yet have converged). Enter θ as a list [θ1,θ2] and θ0 as a single number;

θ = ?

θ0 = ?
Answer #1

θ=[0,0], θo=0

Here, for y=+1, θx+θo>0 and y=-1,θx+θo<=0

if x has label +1 but got misclasfied, then θ=θ+x

if x has label -1 but got misclassified, then θ=θ-x

Using perceptron algorithm,

θ=[-4,2]-2*[-1,-1]-[2,2] =[-4,2]

θo=1-2*1-1=-2

So,θ=[-4,2] and θo=-2